# Belobog: Move Language Fuzzing Framework For Real-World Smart Contracts

### v1

@@ 5.2. RQ2: Coverage and Throughput
We collect the basic-block coverage of each smart contract during fuzzing. Figure 8 shows the averaged results in each smart contract. Across all 67 smart contracts in DS2 and DS1, Belobog reaches significantly higher coverage than ItyFuzz on all of them. Specifically, Belobog reaches an average of 55.6% coverage, while ItyFuzz only reaches 8.8%.
Figure 9 shows the throughput of Belobog and ItyFuzz. Belobog has an average of 129K transaction executions per second, while ItyFuzz has an average of 10K transaction executions per second. Belobog achieves much higher throughput than ItyFuzz in nearly all of the smart contracts, except for the smart contract d46b, on which both fuzzers achieve similar throughput. The core reason is that this smart contract is relatively simple, without many type parameters and “Hot Potato” object types and thus ItyFuzz is easier to find objects of expected types to build seeds. The overall high throughput of Belobog is because we optimized our implementation to incrementally build transactions during fuzzing, avoiding redundant transaction builds. However, ItyFuzz implements a snapshot mechanism, which greatly suffers from the state explosion.
1
2public fun mul_div(a: u128, b: u128, c:u128): u128 {
[…]
4}
5public fun compute_boost_factor(vestar_amount: u128, farm_amount: u128, total: u128): u64 {
6// precision loss because division before multiplication
[…]
8let divisor = mul_div(farm_amount, FACTOR * 2, total);
9((dividend / divisor) + FACTOR) / (FACTOR / 100)
10}
11public fun calculate_boost_weight(amount: u128, boost_factor: u64): u128 {
12// precision loss because division before multiplication
[…]
14}
15public fun update_boost_factor(...) {
16let boost_factor = compute_boost_factor(...);
17let weight = calculate_boost_weight(asset_amount, boost_factor);
18...
19}
Figure 10. A Precision Loss defect of the starswap-core-aptos project.
Table 2. Bug-finding results of Belobog and its variants on DS1. GroundTruth denotes the number of ground-truth bugs.
@@ 5.3. RQ3: Ablation Study
In order to understand the impact of individual components in Belobog, we additionally implement three variants by excluding a few components and evaluate their bug-finding capabilities on DS1.
  • Belobog-NTG: In this variant, we no longer generate the type graph but instead use a random havoc mutation like ItyFuzz.
  • Belobog-NTY: In this variant, we do not fill type parameters.
@@
The last column in Table 2 shows the result of removing all three key components in Belobog. Only two bugs can be detected, suggesting the criticality of the three components. Below we discuss the results of each variant.
Impact of type graph. The Belobog-NTG column in Table 2 shows the bug-finding result after removing the type graph. Compared to our default Belobog, none of the Critical bugs can be detected, and the overall bug detection ratio drops from 85% to nearly zero. The main reason is that without type graphs, Belobog-NTG can no longer guanratee that “Hot Potato” objects are consumed properly and thus many functions can not be covered.
Impact of type parameters. The Belobog-NTY column in Table 2 shows the bug-finding result after removing the support of type parameter. Compared to our default Belobog, only 4 Critical bugs can be detected, and the overall bug detection ratio drops from 85% to 53%. The reason for the bad performance is that 37% of the functions in DS1 have at least one type parameter and thus all transactions generated by Belobog-NTY can not make a valid call to them.
Impact of concolic executor. The Belobog-NCE column in Table 2 shows the bug-finding result after removing the concolic exector. Compared to our default Belobog, only 5 Critical bugs can be detected, and the overall bug detection ratio drops from 85% to 60%. This suggests that concolic execution is crucial to exploring critical paths in smart contracts.
1
2fun check_tick_range(lower: I32, upper: I32, tick_spacing: u32) {
3assert!(lower < upper &&
4lower >= MIN_TICK && upper <= MAX_TICK &&
5lower }
6public fun create_position(ticker_lower: u64, tick_upper: u64, ...): Position {
7check_tick_range(tick_lower, tick_upper, TICK_SPACING);
8..
9}
Figure 11. A check from the smart contract e3564d47.
5.4. Case Study
We provide additional cases from DS1 to demonstrate the capability of Belobog.
Case 1. Figure 11 shows an example smart contract derived from e3564d47 to illustrate how our concolic execution digs into the deep state of a function. The public function create_position() internally will check against the parameters with check_tick_range(), which further creates a rather tight bound on the input value tick_lower and tick_upper. ItyFuzz will never reach the code after the check_tick_range() because the chance of satisfying the constraints by random mutation is negligible, while Belobog can easily pass the assertion with our concolic execution.
Case 2. Compared to MoveScan, Belobog covers every finding it reports with no false positives. We additionally find 27 Precision Loss. We attribute such advantages to two reasons: (1) MoveScan scans these vulnerabilities within a single function context statically, while Belobog is capable of tracking values and invariants across multiple call boundaries at runtime; (2) the concolic execution of Belobog can provide the precise context for oracles like Precision Loss, Infinite Loop and Unnecessary Bool Judgement.
We demonstrate this with two Precision Loss vulnerabilities from the same smart contract from the starswap-core-aptos smart contract of DS2 on the Aptos chain. The first Precision Loss happens at Line LABEL:line:loss1 in Figure 10 because the division of total happens before the multiplication of FACTOR. The second Precision Loss is at Line LABEL:line:loss2 in Figure 10, where the division of FACTOR to boost_factor() happens in the function compute_boost_factor() before the multiplication of amount in calculate_boost_weight() when calling the function update_boost_factor(). MoveScan misses both vulnerabilities because it only detects within a single function call context and does not trace values across function boundaries. In addition, it only supports a very limited fixed pattern like (a / b) * c with exactly three operands. Belobog finds both Precision Loss vulnerabilities because we track symbolic expressions for each value across functions and accurately identify potential precision loss.
5.5. RQ4: Real-world Incident Study
In order to understand the real-world vulnerability detection capability of Belobog, we select two recent catastrophic incidents on the Sui blockchain, known as Cetus and Nemo, which resulted in more than \$200 million and \$2.6 million financial loss, respectively.
@@ 5.5. RQ4: Real-world Incident Study
Cetus Incident Analysis. Figure  12 presents a simplified Cetus smart contract. Although most arithmetic overflows are properly checked by the MoveVM, left shift is not. Thus, there is a left shift check function checked_shlw() implemented in the Cetus to prevent such overflows. However, the function incorrectly encodes the left shift overflow threshold as v > (0xFFFFFFFFFFFFFFFF << 192), which is much bigger than it is supposed to be, i.e., v >= (1 << 192) as the later patch shows. Therefore, any values within the two boundaries, i.e., (1 << 192) < v <= (0xFFFFFFFFFFFFFFFF << 192), will lead to an overflowed value be returned.
@@
with a carefully constructed value AMOUNT. The add_liquidity() calls get_delta() to calculate the tokens (money) that need to be paid back later, and get_delta() calls checked_shlw() to avoid too large values. However, a carefully crafted large value AMOUNT can escape the overflow check in checked_shlw(), leading to a much smaller repay_amout in line LABEL:line:repaya_amount. In other words, the transaction sender can drain a considerable amount of assets with little cost.
Fuzzing Cetus with Belobog. As shown in Figure 12, the two functions, add_liquidity() and repay_liquidity(), have to respectively produce and consume a “Hot Potato” object that has the type AddLituidityReceipt<T>. Therefore, with our built type graph, Belobog is capable of generating the expected function call sequence. However, the type graph itself alone is not enough to trigger this vulnerability for two reasons: (1) triggering the overflow within checked_shlw() is not trivial because the last line of get_delta() casts the value from u256 to u64, resulting in a much tighter bound. Our calculation reveals that the chance of randomly mutating a value to trigger a valid overflow is as small as $2^{-30}$; (2) line LABEL:line:assert_coin_value in repay_liquidity() enforces that the input coin has to be equal to the repay_amout calculated in line LABEL:line:repaya_amount. These constraints make it extremely challenging to generate a successful exploit. Fortunately, Belobog can easily resolve both obstacles with its concolic executor.
When fuzzing starts, Belobog forks the Sui blockchain at checkpoint 148114817, right before the incident happens. Then, it initializes the on-chain object pool with available shared objects obtained from the forked chain. Belobog is capable of detecting the SHL overflow issue in a few seconds and reproducing the whole exploit with the Earning Profits oracle in less than 3 hours, where most of the time is spent on executing transactions.
@@
2fun update_price(v: u64): u64 { ... }
3fun get_price(): u64 { ... }
4fun current_price(price: u64): u64 {
5// Unexpected price updating
6update_price(price);
7get_price()
8}
@@
10amount: u64, price: u64
11): u64 {
12// Trigger the price update
13let price = current_price(price);
14...
@@
(a) The root cause of Nemo.
1
2public struct Position has store { ... }
3public struct FlashPosition<T> { ... }
4public struct Oracle { ... }
@@
6position: Position, oracle: Oracle
7): Position { calculate_amount_by_price(...) }
8public fun borrow<T>(
9position: Position, amount: u64
10): (Position, FlashPosition<T>) { ... }
11public func repay<T>(
12position: Position, receipt: FlashPosition<T>
13) { ... }
14public func get_oracle(...): Oracle { ... }
(b) Simplified smart contract of Nemo where exploits interacts.
Figure 13. Simplified source code of Nemo.
Nemo Incident Analysis. The root cause of Nemo is an access control issue as presented in the Figure 13(a). The function calculate_amount_by_price() is designed to estimate the output amount by the given amount and price. However, its underlying private function current_price() writes the user input price to its internal states. Therefore, anyone can potentially earn profits by calling the function to manipulate the price. Note that although triggering the vulnerability in Figure 13(a) is straightforward, it is not trivial to generate a full exploit because modifying the price does not yield profits directly. From a very high level, an attacker needs the following steps interacting with the functions, which is shown in Figure 13(b). A successfull exploit should: (1) borrow a flash loan via borrow<T>() to get some assets T0, (2) call several swap<T0,T1>() and get_oracle() functions to swap the assets from T0 to T1; (3) manipulate the price of T0 and T1 with calculate_amount_by_price(); (4) swap T1 back to T0 with the shifted price; and (5) pay back the flashloan via repay<T>(). In practice, the actual exploit (nemoexp) composes a sequence of 231 Move function calls.
Fuzzing Nemo with Belobog. Thanks to the type graph, Belobog can always ensure the correct pairs of swap<T0,T1>() with get_oracle() and borrow<T>() with repay<T>(). This greatly reduces the sequence search space and enables Belobog to efficiently assemble the exploit with 10 Move function calls to trigger the Earning Profits oracle. Belobog forks the Sui blockchain at checkpoint 180415392, right before the Nemo attack, and fills in the related shared objects. It can generate a whole profitable exploit in less than 30 minutes.
@@ 6. Discussion
Automated Testing Smart Contracts. Both Cetus and Nemo studied in our evaluation are heavily audited by the leading Web3 auditing companies, but still, they fail to prevent these critical vulnerabilities. This suggests that humans are prone to making errors, and pure manual efforts are not enough to protect the security of smart contracts. Therefore, continuous and automated testing for the smart contracts is essential to secure deployed smart contracts. While there are already lots of automated testing tools on the prevailing Ethereum Virtual Machine (EVM), we find that the Move ecosystem greatly lacks such tools and frameworks. Therefore, we hope that Belobog can shed light on the Move smart contracts testing by providing an extensible and general fuzz testing framework.
Oracle Automation. While Belobog has demonstrated its effectiveness in finding bugs and achieved high coverage and throughput, one of the limitations is that Belobog needs extra manual efforts to build oracles when testing real-world smart contracts that have a custom business model. We workaround this by providing a wide range of predefined oracles for common defects and allowing users to write oracles in Move language instead of modifying the fuzzer itself. It is an interesting direction to explore the automated fuzzing oracle synthesis for the Move smart contracts, given their strongly typed nature. It is possible to apply some existing approaches (actlifter; defiranger; defort; sleuth) to extract high-level smart contract semantics to help build oracles automatically. We deem this as one of the interesting directions for our future research.
State Dependency. Although Belobog builds a type graph to model the type dependencies of various functions, it does not model the inner state dependency. In practice, smart contracts are stateful programs on the blockchain. Our evaluation suggests that our concolic executor could solve the direct dependency bound to the input parameters as the Cetus study in Section 5.5. It is also possible to integrate dataflow analysis to understand the state dependency, like Smartian(smartian). We leave this as our future work.
@@ 7. Related Work
Move Language Testing. So far, there is a limited number of tools to test Move contracts automatically. Move Prover (moveprover) pioneers in translating the Move language to its custom intermediate verification language to verify pre- and post-conditions. However, it requires manual effort to write specifications and only supports a limited set of oracles, making it less practical given the fancy business models of the DeFi ecosystem and learning costs (movescan). MoveLint (movelint) is an open-sourced tool that detects several Move defects by analyzing the Abstract Syntax Tree of smart contracts. While it supports a wide range of oracles, it suffers from false positives due to the inaccuracy of static rules. MoveScan (movescan) instead translates the Move bytecode into stackless intermediate representation and generates Control Flow Graph to facilitate code analysis with extended oracle definitions. ItyFuzz (ityfuzz) is originally designed for fuzzing bytecode on Ethereum Virtual Machine (EVM), it also ports its snapshot-based fuzzing approaches to MoveVM afterwards. Sui-Fuzzer (suifuzzer) is another coverage-guided fuzzer supporting both stateful and stateless fuzzing with custom properties.
@@
Smart Contract Fuzzing. Fuzzing is an efficient way to test smart contracts. sFuzz (sfuzz) uses strategies adapted from AFL to improve the code coverage. Smartian (smartian) proposes both static and dynamic dataflow analysis to solve the complex data dependencies. ConFuzzius (confuzzius) hybrids symbolic execution, data dependency analysis, and taint analysis to exercise deeper bugs. VULSEYE (vulseye) novelly models the smart contract fuzzing as stateful directed fuzzing and guides the fuzzing by vulnerabilities. VERITE (verite) builds a first profit-centric fuzing framework with a gradient descent method to maximize the exploits. LLM4Fuzz (llm4fuzz) attempts to utilize the capability of Large Language Model (LLM) to guide the smart contract fuzzing. However, all of them focus on smart contracts on EVM and do not consider various type constraints in smart contracts. Thus, it is non-trivial to port them to support Move smart contracts as we have tried on ItyFuzz.
@@
API Fuzzing. Generating a sequence of function calls to test the smart contract is similar to the traditional API fuzzing for libraries. GraphFuzz (graphfuzzing) models the execution trace as a dataflow graph to test the low-level library APIs. Minerva (minerva) utilizes dynamic mod-ref analysis to test the browser APIs. APICraft (apicraft) leverages both static and dynamic information to build fuzzer driver for testing closed-source SDKs. RESTler (restler) as a stateful fuzzer analyzes the specification of the REST API by inferring the producer and consumer relations like our type graph building process. OGHARN (noharness) proposes the novel oracle-guided harnessing to automatically fuzz API. Due to the special requirements and constraints of smart contracts, applying such general API fuzzers to smart contracts cannot work as it is. However, it is interesting to explore how various API fuzzing insights can be applied in smart contract fuzzing.
8. Conclustion

### v2

@@ 5.2. RQ2: Coverage and Throughput
We collect the edge coverage of each project during fuzzing. Figure 7 shows the average results in each project. Across all 36 out of 37 projects in DS1, Belobog reaches significantly higher coverage than ItyFuzz on all of them. Specifically, Belobog achieves an average coverage of 55.6%, whereas ItyFuzz achieves only 8.8%. ItyFuzz reaches similar code coverage to Belobog only on project 73c0. The reason is that the project needs administrator privileges for deeper testing, which neither Belobog nor ItyFuzz assumes. For the fuzzing throughput, Belobog is ten times faster than ItyFuzz. The reason is that ItyFuzz is type unaware, leading to lots of computation waste in enumerating objects and calls to consume the hot potatoes. We further include a detailed throughput evaluation in our supplementary materials.
Table 3. Bug-reachability results of Belobog and its variants on DS1.
@@ 5.3. RQ3: Ablation Study
In order to understand the impact of individual components in Belobog, we additionally implement four variants by excluding core components and evaluate their bug-reachability capabilities on DS1.
@@
  • Belobog-NTG: In this variant, we no longer generate the type graph but instead use a random havoc mutator as implemented in ItyFuzz.
  • Belobog-NTY: In this variant, we do not fill type parameters and skip functions requiring type parameters.
@@
The last column in Table 3 shows the result of removing all three key components in Belobog. Only two bugs can be detected, suggesting the criticality of the three components. Below, we discuss the results of each variant.
Impact of type graph. The Belobog-NTG column in Table 3 shows the bug-reachability result after removing the type graph. Compared to our default Belobog, none of the Critical bugs can be detected, and the overall bug detection ratio drops from 85% to nearly zero. The main reason is that without type graphs, Belobog-NTG can no longer guarantee that “Hot Potato” objects are consumed properly, and thus many functions can not be covered.
Impact of type parameters. The Belobog-NTY column in Table 3 shows the bug-reachability result after removing the support of type parameter. Compared to our default Belobog, only 4 Critical bugs can be detected, and the overall bug detection ratio drops from 85% to 53%. The reason for the bad performance is that 37% of the functions in DS1 have at least one type parameter and thus all transactions generated by Belobog-NTY can not make a valid call to them.
Impact of concolic executor. The Belobog-NCE column in Table 3 shows the bug-reachability result after removing the concolic executor. Compared to our default Belobog, only 5 Critical bugs can be detected, and the overall bug detection ratio drops from 85% to 60%. This suggests that concolic execution is crucial to exploring critical paths in smart contracts.
@@
5.4. RQ4: On-Chain Incidents Reproduction
We select two recent catastrophic on-chain incidents on the Sui blockchain, known as Cetus (3) and Nemo (18), which resulted in more than \$200 million and \$2.6 million financial losses, respectively. Belobog is capable of reproducing the full exploits of both incidents with just the Earning Profits oracle. We discuss Cetus in this section and Nemo in our supplementary material.
Figure  8 presents a simplified Cetus smart contract. Although most arithmetic overflows are properly checked by the MoveVM, left shift is not. Thus, there is a left shift check function checked_shlw() implemented in the Cetus to prevent such overflows. However, the function incorrectly encodes the left shift overflow threshold as v > (0xFFFFFFFFFFFFFFFF << 192) (Line LABEL:line:leftshift), which is much bigger than it is supposed to be, i.e., v >= (1 << 192) as the later patch shows. Therefore, any values between the two boundaries, i.e., (1 << 192) < v <= (0xFFFFFFFFFFFFFFFF << 192), will lead to an overflowed value being returned.
@@ 5.4. RQ4: On-Chain Incidents Reproduction
with a carefully constructed value AMOUNT. The add_liquidity() calls get_delta() to calculate the tokens (money) that need to be paid back later, and get_delta() calls checked_shlw() to avoid too large values. However, a carefully crafted large value AMOUNT can escape the overflow check in checked_shlw(), leading to a much smaller repay_amount in line LABEL:line:repaya_amount. In other words, the transaction sender can drain a considerable amount of assets with little cost.
Fuzzing Cetus with Belobog. As shown in Figure 8, the two functions, add_liquidity() and repay_liquidity(), have to respectively produce and consume a “Hot Potato” object that has the type AddLiquidityReceipt<T>. Therefore, using our built-type graph, Belobog can generate the expected function-call sequence. However, the type graph itself alone is not enough to trigger this vulnerability for two reasons: (1) triggering the overflow within checked_shlw() is not trivial because the last line of get_delta() casts the value from u256 to u64, resulting in a much tighter bound. Our calculation reveals that the chance of randomly mutating a value to trigger a valid overflow is as small as $2^{-30}$; (2) line LABEL:line:assert_coin_value in repay_liquidity() enforces that the input coin has to be equal to the repay_amount calculated in line LABEL:line:repaya_amount. These constraints make it extremely challenging to generate a successful exploit. Fortunately, Belobog can readily address both obstacles using its concolic executor.
When fuzzing starts, Belobog forks the Sui blockchain at checkpoint 148114817, right before the incident happens. Then, it initializes the on-chain object pool with available shared objects obtained from the forked chain. Belobog is capable of detecting the SHL overflow issue in a few seconds and reproducing the whole exploit with the Earning Profits oracle in less than 3 hours, where most of the time is spent on executing seeds.
5.5. RQ5: Real-world Application
We further apply Belobog on another three projects that are still under auditing. Belobog can exploit one critical vulnerability leading to draining pool rewards by Earning Profits and one medium vulnerability by Precision Loss without manually crafting oracles. By further implementing oracles for the projects, Belobog additionally reports 1 critical, 2 major, and 2 medium vulnerabilities. All these findings have been confirmed by the project developers and fixed before deployment. We will showcase how Belobog exploits the critical vulnerability to earn profits while leaving other vulnerabilities in our supplementary materials.
@@ 5.5. RQ5: Real-world Application
3public fun open_position<X, Y>(...): Position {
4Position { liquidity: 0, fee_growth_last: 0, rewards: 0, }
5}
6public fun add_liquidity<X, Y>(position, liquidity, ...) { position.update(liquidity); }
7/// internal helper
8fun update(position, delta, ...){
[…]
10// rewards are non-empty for newly created positions because position.liquidity is not zero
[…]
12position.rewards += rewards; // position.rewards should be zero for newly created positions
13position.fee_growth_last = pool_fee_growth;
14// Fix: update position.liquidity here
@@
Figure 9. A simplified code snippet of a real-world bug that allows anyone to drain pool rewards.
Real-world Bug Incident Study. The simplified code of the target project, as illustrated in the Figure 9, is a Decentralized Exchange (DEX) that accumulates fees into pool_fee_growth (Line LABEL:line:fee_growth) that are later distributed to the positions of liquidity providers (Line LABEL:line:distribute_fee) as rewards.
However, the code that updates position liquidity has a serious flaw: it calculates rewards based on the updated position liquidity as Line LABEL:line:update_bug suggests, which causes a newly created position to receive non-zero rewards from position.rewards. This allows any user to withdraw all previous rewards by creating a position with the minimum liquidity requirement using the sequence
$\texttt{open\_position()}\Rightarrow\texttt{add\_liquidity(position, 1)}\Rightarrow\texttt{collect\_fees(position)}$
Fuzzing with Belobog Belobog will compile the target package and deploy it on a locally forked blockchain environment. Belobog can generate the full exploit sequence in less than 30 minutes because it triggers the Earning Profits oracle by withdrawing the rewards of previous swaps.
@@ 6. Discussion
Automated Testing Smart Contracts. Both Cetus and Nemo studied in our evaluation are heavily audited by the leading Web3 auditing companies, but still, they fail to prevent these critical vulnerabilities. This suggests that humans are prone to making errors, and pure manual efforts are not enough to protect the security of smart contracts. With Belobog, these losses could have been avoided. Therefore, continuous and automated testing for the smart contracts is essential to secure deployed smart contracts. While there are already lots of automated testing tools on the prevailing Ethereum Virtual Machine (EVM), we find that the Move ecosystem greatly lacks such tools and frameworks. Therefore, we hope that Belobog can shed light on the Move smart contracts testing by providing an extensible and general fuzz testing framework.
Oracle Automation. While Belobog has demonstrated its effectiveness in finding bugs and achieved high coverage and throughput, one of the limitations is that Belobog needs extra manual efforts to build oracles when testing real-world smart contracts that have a custom business model. We work around this by providing a wide range of predefined oracles for common defects and allowing users to write oracles in Move language instead of modifying the fuzzer itself. It is an interesting direction to explore the automated fuzzing oracle synthesis for the Move smart contracts, given their strongly typed nature. It is possible to apply some existing approaches (Li et al., 2023; Wu et al., 2023; Xie et al., 2024; Wu et al., 2024) to extract high-level smart contract semantics to help build oracles automatically. We deem this as one of the interesting directions for our future research.
State Dependency. Although Belobog builds a type graph to model the type dependencies of various functions, it does not model the inner state dependency. In practice, smart contracts are stateful programs on the blockchain. Our evaluation suggests that our concolic executor could solve the direct dependency bound to the input parameters as the Cetus study in Section 5.4. It is also possible to integrate dataflow analysis to understand the state dependency, like Smartian(Choi et al., 2021). We leave this as our future work.
@@ 7. Related Work
Move Language Testing. So far, there are a limited number of tools to test Move contracts automatically. Move Prover (Zhong et al., 2020) is a pioneer in translating the Move language to its custom intermediate verification language to verify pre- and post-conditions. However, it requires manual effort to write specifications and only supports a limited set of oracles, making it less practical given the fancy business models of the DeFi ecosystem and learning costs (Song et al., 2024). MoveLint (16) is an open-source tool that detects several Move defects by analyzing the Abstract Syntax Tree of smart contracts. While it supports a wide range of oracles, it suffers from false positives due to the inaccuracy of static rules. MoveScan (Song et al., 2024) instead translates the Move bytecode into stackless intermediate representation and generates a Control Flow Graph to facilitate code analysis with extended oracle definitions. ItyFuzz (Shou et al., 2023) is originally designed for fuzzing bytecode on Ethereum Virtual Machine (EVM), it also ports its snapshot-based fuzzing approaches to MoveVM afterwards. Sui-Fuzzer (27) is an coverage-guided fuzzer supporting both stateful and stateless fuzzing with custom properties.
@@
Smart Contract Fuzzing. Fuzzing is an efficient way to test smart contracts. sFuzz (Nguyen et al., 2020) uses strategies adapted from AFL to improve the code coverage. Smartian (Choi et al., 2021) proposes both static and dynamic dataflow analysis to solve the complex data dependencies. ConFuzzius (Torres et al., 2021) hybridizes symbolic execution, data dependency analysis, and taint analysis to exercise deeper bugs. VULSEYE (Liang et al., 2025) novelly models the smart contract fuzzing as stateful directed fuzzing and guides the fuzzing by vulnerabilities. VERITE (Kong et al., 2025) builds a first profit-centric fuzzing framework with a gradient descent method to maximize the exploits. LLM4Fuzz (Shou et al., 2024) attempts to utilize the capability of Large Language Model (LLM) to guide the smart contract fuzzing. However, all of them focus on smart contracts on EVM and do not consider various type constraints in smart contracts. Thus, it is non-trivial to port them to support Move smart contracts as we have tried with ItyFuzz. Meanwhile, it would also be hard to port our methods to Solidity and EVM because of its hashing table style storage, where type information is much less visible from the public function signatures. For instance, the interface of standard ERC20 (5) that represents assets on EVM uses primitive values like address and uint256 while Move has strongly-typed Coin¡T¿ for assets. Therefore, the benefits of building type graphs for public interfaces on Solidity and EVM are very limited.
@@
API Fuzzing. Generating a sequence of function calls to test the smart contract is similar to the traditional API fuzzing for libraries. GraphFuzz (Green and Avgerinos, 2022) models the execution trace as a dataflow graph to test the low-level library APIs. Minerva (Zhou et al., 2022) utilizes dynamic mod-ref analysis to test the browser APIs. APICraft (Zhang et al., 2021) leverages both static and dynamic information to build a fuzzer driver for testing closed-source SDKs. OGHARN (Sherman and Nagy, 2025) proposes the novel oracle-guided harnessing to automatically fuzz APIs. There are also several works utilizing type information to test code, like RPG (Xu et al., 2024), RuMono (Zhang et al., 2025), TypeNFuzz (Zeng et al., ), and SYPET (Feng et al., 2017). However, the Move Language type safety rules, including the “Hot Potato” mechanism, present additional challenges and gaps. Therefore, applying such methods can not work as it is, though it is interesting to explore how these insights could help enhance the capabilities of Belobog.
8. Conclusion